OpenStudy (anonymous):

The voltage and current at the terminals of this circuit are zero for t < 0 and for t ≥ 0 they are v = 54 e(^-1573t) Volts i = 5.6e(^-1573t) Amps

OpenStudy (anonymous):

Find power and energy delivered at t = .000625 seconds in watts and milliJoules

OpenStudy (anonymous):

My calculations gave me: P(625microseconds) = 36.332 watts Work = 69 milliJoules

OpenStudy (irishboy123):



OpenStudy (irishboy123):

you are using \(V = 54 e^{-1573t} \) and \( I = 5.6e^{-1573t}\) so \(P(t) = 54(5.6)e^{-2 \times 1573t}\) and \(W(t) = \int\limits_{0}^{0.000625} P(t) \ dt = \left[ \frac{54(5.6)}{-2 \times 1573}e^{-2 \times 1573t} \right]_{0}^{0.000625}\) is that your approach? because i'm getting different answers

OpenStudy (anonymous):

Apologies late reply, so late... But you're approach is the correct approach. My mistake was that I took the instantaneous value of one of my variables before plugging it into the Work Integral.